# Turing’s Test, a Beautiful Thought Experiment

### v2

@@ 7. The value of Turing’s test for AI
By May 1953, John McCarthy and Claude Shannon were working on their collection Automata Studies (McCarthy and Shannon, 1956), which revolved around “the theory of Turing machines” (p. vii), and to which they invited Turing to contribute. [^30] Shannon and McCarthy to Turing, May 18, 1953. Alan Turing Papers (Additional), University of Manchester Library, GB133 TUR/Add/123. Turing declined the invitation, saying that he had been working for the last two years on “the mathematics of morphogenesis,” although he expected “to get back to cybernetics very shortly.” [^31] Turing to Shannon, June 3, 1953 (ibid.). One year and four days later, Turing was dead, and early AI would not note his biological turn. Commenting on “the Turing definition of thinking” (p. vi), McCarthy and Shannon found it “interesting” because it “has the advantages of being operational or, in the psychologists’ term, behavioristic $\ldots$ No metaphysical notions of consciousness, ego and the like are involved.” They also thought that this very strength could be a weakness, because it has “the disadvantage” of being susceptible to a memorizing machine playing the imitation game by looking up “a suitable dictionary.”
@@
McCarthy and Shannon referred interchangeably to ‘definition’ and to a word that Turing actually used, ‘criterion:’ “While certainly no machines at the present time can even make a start at satisfying this rather strong criterion, Turing has speculated that within a few decades it will be possible to program general purpose computers in such a way as to satisfy this test” (McCarthy and Shannon, 1956, p. v, emphasis added).
@@
In 1955, before the publication of Automata Studies, McCarthy and Shannon, together with Marvin Minsky and Nathaniel Rochester, co-authored their well-known ‘Proposal’ for AI research (McCarthy et al., 1955). Unlike Turing himself, they seem to have thought of machine intelligence in terms of Turing machines, as their opening paragraph suggests: “The study is to proceed on the basis of the conjecture that every aspect of learning or any other feature of intelligence can in principle be so precisely described that a machine can be made to simulate it.” Whether or not they followed Turing on the ‘how,’ they certainly followed him on the ‘what,’ in writing: “For the present purpose the artificial intelligence problem is taken to be that of making a machine behave in ways that would be called intelligent if a human were so behaving” (p. 7). This definition — compare it with “the Turing definition of thinking” — would stay.
@@
In the early 1960s, Edward Feigenbaum and Julian Feldman noted in Computers and Thought (Feigenbaum and Feldman, 1963) that Turing’s 1950 paper “appeared five years before concrete developments in intelligent behavior by machine began to occur;” and “yet,” they continued, “it remains today one of the most cogent and thorough discussions in the literature on the general question “Can a machine think?” (pp. 9-10). They observed Turing’s “behavioristic posture relative to the question,” which “is to be decided by an unprejudiced comparison of the alleged ‘thinking behavior’ of the machine with normal ‘thinking behavior’ in human beings” (emphasis added). They concluded: “He proposes an experiment — commonly called ‘Turing’s test’ — in which the unprejudiced comparison could be made $\ldots$ Though the test has flaws, it is the best that has been proposed to date.”
@@
Minsky, in the preface to his 1967 collection (Minsky, 1968), reiterates the definition of AI as “the science of making machines do things that would require intelligence if done by men” (p. v). Around the same time, Minsky collaborated with Stanley Kubrick and Arthur Clarke on their 1968 screenplay, also written as a novel, 2001: A Space Odyssey (Clarke, 1968), which featured a futuristic computer named HAL:
@@
The “Turing definition of thinking” was to become legendary.
@@
Stuart Shieber studied McCarthy and Shannon’s memorizing machine objection in depth, elaborated on its assumptions, and concluded that it is invalid (Shieber, 2014). But McCarthy’s concept of ‘memorizing’ may have been more elastic, as his later comment on Deep Blue’s defeat of Gary Kasparov seems to suggest (McCarthy, 1997). He expressed disappointment that it was mostly an achievement of computational power rather than thinking, and gave a clear argument why he thought so. Essentially, McCarthy pointed out, computer chess advanced by replacing heuristic techniques, which relied on the expertise of human players to prune the search space of possible moves, with brute force computing. “[I]t is a measure of our limited understanding of the principles of artificial intelligence,” McCarthy wrote, “that this level of play requires many millions of times as much computing as a human chess player does.” It may be, but that the problem was “largely a quantitative matter” was hinted at by Turing in his letter to Worsley of c. June 1951 (Box 3).
@@
Ten years after Deep Blue vs. Kasparov, McCarthy referred to Turing’s 1947 lecture (op. cit.) as “the first scientific discussion of human level machine intelligence,” and to Turing’s 1950 paper as “amplifying” that discussion into a “goal” (McCarthy, 2007, p. 1174).
@@
In 1992, Minsky co-authored a work of fiction, The Turing Option (Warner, New York), in which Turing’s test is featured in the preface. In 1995, Minsky took a stand against Loebner’s Weinzenbaum experiments, pleading to “revoke his stupid prize, save himself some money, and spare us the horror of this obnoxious and unproductive annual publicity campaign.”323232‘Annual Minsky Loebner Prize Revocation Prize 1995 Announcement,’ 2 March 1995. Available at: https://groups.google.com/g/comp.ai/c/dZtU8vDD_bk/m/QYaYB18qAToJ. Accessed 25 Nov 2023. In 2013, when asked about the Turing test in a taped interview, Minsky said: “The Turing test is a joke, sort of, about saying ‘A machine would be intelligent if it does things that an observer would say must be being done by a human’ $\ldots$ it was suggested by Alan Turing as one way to evaluate a machine but he had never intended it as being the way to decide whether a machine was really intelligent.”333333‘Marvin Minsky on AI: the Turing test is a joke!’, from 23’ 35” to 24’45”. Available at https://www.singularityweblog.com/marvin-minsky/. Accessed Dec. 6, 2023. This materially connects McCarthy et al.’s definition of “the AI problem” with Turing’s test, if material evidence were still needed.
@@
Overall, it seems that all of these AI pioneers understood and were inspired by Turing’s test at the level of conceptual foundations. Even if some of them also used the term ‘experiment,’ none of them took it literally as a practical experiment, which would indeed imply an astonishing lack of imagination on their part. Turing’s test moved the burgeoning field of AI away from unproductive debates about the meaning of words, for example, allowing Minsky to write in 1967 (Minsky, 1967): “Turing discusses some of these issues in his brilliant article, ‘Comput­ing Machines and Intelligence” [sic], and I will not recapitulate his arguments $\ldots$ They amount, in my view, to a satisfactory refutation of many such objections” (p. 107).
@@
The value of Turing’s test ($P_{5}$) is that it has long been and still is a unifying ‘definition,’ a ‘criterion,’ a ‘goal’ for, in the words of McCarthy et al., the science and engineering of “making a machine behave in ways that would be called intelligent if a human were so behaving.” Every time AI succeeds in automating a new task that was once reserved for humans because it requires intelligence, “the Turing definition” conquers new territory, and the significance of Turing’s early message to his contemporaries becomes clearer.
8. Conclusion
In this paper, I presented a mass of evidence, including newly discovered archival sources, and a new perspective on Turing’s test.
@@ 8. Conclusion
New light was shed on Turing’s concept of imitation, emphasizing that it does not encourage deception in AI. Rather, it is a mathematical concept, largely in continuity with his 1936 paper. I also showed that Turing’s presentation of the various versions of his test fits what Mach characterized as “the basic method of thought experiments” in science. I reconstructed the historical conditions of Turing’s proposal, explaining that gender imitation was his ironic response to Jefferson, and conversation was yet another concession to his opponents. Further, I compared Turing’s test to Galileo’s falling-bodies experiment, showing that the problem of circularity is inherent in existential hypotheses, and that the solution may lie in propaganda and the progressive scientific and technological developments of a next generation of scientists. I then revisited the history of AI, showing that Turing’s test provided McCarthy, Minsky, and others with a definition of the AI problem that, at the level of conceptual foundations, still drives AI research today.
But whatever its utility, Turing’s test has secured its place as one of the most beautiful thought experiments in the history of science.
Acknowledgements.
@@
The author thanks Andrew Hodges, Jim Miles, and H. V. Jagadish for their valuable comments on an earlier version of this article, Mark Priestley for the gift of Turing’s letters to Worsley, and Fabio Cozman and Murray Shanahan for their support. The author is solely responsible for the accuracy of this work. This research has been supported by the São Paulo Research Foundation (FAPESP grants 2022/16793-9 and 2019/21489-4, “The Future of Artificial Intelligence: The Logical Structure of Alan Turing’s Argument”), and by the IBM Corporation and the São Paulo Research Foundation (FAPESP grant 2019/07665-4).

### v3

@@ Abstract
By May 1953, Shannon was working with John McCarthy on their collection Automata Studies [mccarthy1956], which revolved around “the theory of Turing machines” (p. vii), and to which they invited Turing to contribute. [^36] Shannon and McCarthy to Turing, May 18, 1953. Alan Turing Papers (Additional), University of Manchester Library, GB133 TUR/Add/123. Turing declined the invitation, saying that he had been working for the last two years on “the mathematics of morphogenesis,” although he expected “to get back to cybernetics very shortly.” [^37] Turing to Shannon, June 3, 1953 (ibid.). One year and four days later, Turing was dead, and early AI would not note his biological turn. Commenting on “the Turing definition of thinking” (p. vi), McCarthy and Shannon found it “interesting” because it “has the advantages of being operational or, in the psychologists’ term, behavioristic $\ldots$ No metaphysical notions of consciousness, ego and the like are involved.” They also thought that this very strength could be a weakness, because it has “the disadvantage” of being susceptible to a memorizing machine playing the imitation game by looking up “a suitable dictionary.”
@@
McCarthy and Shannon referred interchangeably to “definition” and to a word that Turing actually used, “criterion:” “While certainly no machines at the present time can even make a start at satisfying this rather strong criterion, Turing has speculated that within a few decades it will be possible to program general purpose computers in such a way as to satisfy this test” [mccarthy1956, p. v, emphasis added].
@@
In 1955, before the publication of Automata Studies, McCarthy and Shannon, together with Marvin Minsky and Nathaniel Rochester, co-authored their well-known “Proposal” for AI research [mccarthy1955]. Unlike Turing himself, they seem to have thought of machine intelligence in complete continuity with Turing machines, as their opening paragraph suggests: “The study is to proceed on the basis of the conjecture that every aspect of learning or any other feature of intelligence can in principle be so precisely described that a machine can be made to simulate it.” Turing’s postwar view, which we have seen in connection with his concepts of machine and imitation above, seems instead to be that machines would learn their behavior primarily from experience, growing in intelligence like a human child, not always by being given precise instructions on how to do things. In any case, the proponents of the Dartmouth seminar did follow Turing closely in writing: “For the present purpose the artificial intelligence problem is taken to be that of making a machine behave in ways that would be called intelligent if a human were so behaving” (p. 7). This definition — compare it with “the Turing definition of thinking” — would stay. Intelligent machines would be machines designed to match or exceed human performance in a range of cognitive tasks. The implication, emphasized by both Turing and Norbert Wiener but not by the founders of AI, was that humans could no longer be needed for most white-collar jobs in the labor market.
@@
In the early 1960s, E. Feigenbaum and J. Feldman noted in Computers and Thought [feigenbaum1963] that Turing’s 1950 paper “appeared five years before concrete developments in intelligent behavior by machine began to occur;” and “yet,” they continued, “it remains today one of the most cogent and thorough discussions in the literature on the general question “Can a machine think?” (pp. 9-10). They observed Turing’s “behavioristic posture relative to the question,” which “is to be decided by an unprejudiced comparison of the alleged ‘thinking behavior’ of the machine with normal ‘thinking behavior’ in human beings” (emphasis added). They concluded: “He proposes an experiment — commonly called ‘Turing’s test’ — in which the unprejudiced comparison could be made $\ldots$ Though the test has flaws, it is the best that has been proposed to date.”
@@
Minsky, in the preface to his 1967 collection [minsky1968], reiterates the definition of AI as “the science of making machines do things that would require intelligence if done by men” (p. v). Around the same time, Minsky collaborated with Stanley Kubrick and Arthur Clarke on their 1968 screenplay, also written as a novel, 2001: A Space Odyssey [clarke1968], which featured a futuristic computer named HAL:
@@
The “Turing definition of thinking" was to become legendary.
[…]
@@
yellow McCarthy and Shannon’s memorizing machine objection was studied in depth by Stuart Shieber, who elaborated on its assumptions and concluded that it is invalid [shieber2014]. But McCarthy’s concept of memorizing may have been more elastic, as his later comment on Deep Blue’s defeat of Gary Kasparov suggests [mccarthy1997]. He expressed disappointment that it was mostly an achievement of computational power rather than thinking, and gave a clear argument why he thought so. Essentially, computer chess advanced by replacing heuristic techniques, which relied on the expertise of human players to prune the search space of possible moves, with brute force computing. “[I]t is a measure of our limited understanding of the principles of artificial intelligence,” McCarthy wrote, “that this level of play requires many millions of times as much computing as a human chess player does.” It may be, but that the problem was “largely a quantitative matter” was hinted at by Turing in his letter of c. June 1951 (Box 3).
@@
Ten years after Deep Blue vs. Kasparov, McCarthy referred to Turing’s 1947 lecture (op. cit.) as “the first scientific discussion of human level machine intelligence,” and to Turing’s 1950 paper as “amplifying” that discussion into a “goal” [mccarthy2007, p. 1174].
@@
In 1992, Minsky co-authored a work of fiction, The Turing Option (Warner, New York), in which Turing’s test is featured in the preface. In 1995, Minsky took a stand against Loebner’s Weizenbaum experiments, pleading to “revoke his stupid prize, save himself some money, and spare us the horror of this obnoxious and unproductive annual publicity campaign.”383838‘Annual Minsky Loebner Prize Revocation Prize 1995 Announcement,’ 2 March 1995. Available at: https://groups.google.com/g/comp.ai/c/dZtU8vDD_bk/m/QYaYB18qAToJ. Accessed 25 Nov 2023. In 2013, when asked about the Turing test in a taped interview, Minsky said: “The Turing test is a joke, sort of, about saying ‘A machine would be intelligent if it does things that an observer would say must be being done by a human’ $\ldots$ it was suggested by Alan Turing as one way to evaluate a machine but he had never intended it as being the way to decide whether a machine was really intelligent.”393939‘Marvin Minsky on AI: the Turing test is a joke!’, from 23’ 35” to 24’45”. Available at https://www.singularityweblog.com/marvin-minsky/. Accessed Jun. 10, 2024. This materially connects McCarthy et al.’s definition of “the AI problem” with Turing’s test, if material evidence were still needed.
@@
Overall, it seems that all of these AI pioneers understood and were inspired by Turing’s test at the level of conceptual foundations. Even if some of them also used the term “experiment,” none of them took it literally as a practical experiment, which would indeed imply an astonishing lack of imagination on their part. The Turing test helped move the burgeoning field of AI away from unproductive debates about the meaning of words, allowing Minsky, for example, to write in 1967 [minsky1967]: “Turing discusses some of these issues in his brilliant article, ‘Computing Machines and Intelligence’ [sic], and I will not recapitulate his arguments $\ldots$ They amount, in my view, to a satisfactory refutation of many such objections” (p. 107).
@@
The value of Turing’s test is that it has long been and still is a unifying “definition,” a “criterion,” a “goal” for, in the words of McCarthy et al., the science and engineering of “making a machine behave in ways that would be called intelligent if a human were so behaving.” For better or worse, every time AI succeeds in automating a new task that was once reserved for humans because it requires intelligence, “the Turing definition of thinking” conquers new territory, and the significance of Turing’s early message to his contemporaries becomes clearer.
CONCLUSION
@@
This paper presented a new perspective on Turing’s test. New light has been shed on Turing’s concept of imitation, suggesting that it does not give a license for deception in AI. Rather, imitation was for Turing a mathematical concept, largely in continuity with his 1936 paper, although he later generalized how it could be achieved. It was also suggested that Turing’s presentation of the various versions of his test fits what Mach called “the basic method of thought experiments” in the history of science. The historical conditions of Turing’s proposal were reconstructed, showing that the basic idea of a gender test had been raised originally by Jefferson, and Turing’s imitation game comes out of that context. Conversational performance was also a concession to his opponents, and overall Turing’s test was a response to critics. But Turing also took the opportunity to promote his positive views. The known primary and secondary sources indicate that he became actively engaged in machine intelligence propaganda, and it was also in this spirit that he proposed his test, hoping to influence contemporaries and future generations of scientists. The question of the value of Turing’s test and its relation to early AI was revisited, arguing that “the Turing definition of thinking” provided McCarthy, Minsky, and others with a definition of the AI problem at the level of conceptual foundations that arguably still drives AI research today.
But whatever its utility, we can now appreciate that there is more to the imitation game. With its structural elements neatly designed as lighthearted concessions to opponents, and at the same time able to demonstrate the power of digital computing as early as 1950, Turing’s test has secured its place as one of the most beautiful thought experiments in the history of science.
ACKNOWLEDGMENTS
@@
The author thanks Andrew Hodges, Jim Miles, and H. V. Jagadish for their valuable comments on an earlier version of this article; Mark Priestley for the gift of the Turing letters to Worsley; Fabio Cozman and Murray Shanahan for their support; The author is solely responsible for the accuracy of this work. The author thanks the Center for Artificial Intelligence (C4AI-USP) and the support from the São Paulo Research Foundation (FAPESP grants nos. 2019/07665-4, 2019/21489-4, and 2022/16793-9) and from the IBM Corporation. This article is a result of the project “The Future of Artificial Intelligence: The Logical Structure of Alan Turing’s Argument”).
[…]
Bernardo Gonçalves is currently a researcher at the Center for Artificial Intelligence (C4AI), University of São Paulo, Brazil, and a Visiting Fellow at King’s College, University of Cambridge, UK. He works on Alan Turing, AI and computer science. He received Ph.D. degrees in Philosophy from the University of São Paulo and in Computational Modeling from the National Laboratory for Scientific Computing, Brazil.